# Impact of the continuum Coulomb interaction in quantum-orbit-based treatments of high-order above-threshold ionization

### 08 Dec 2023 v1

@@ Abstract
We perform a systematic comparison between photoelectron momentum distributions computed with the rescattered-quantum orbit strong-field approximation (RQSFA) and the Coulomb-quantum orbit strong-field approximation (CQSFA). We exclude direct, hybrid, and multiple scattered CQSFA trajectories, and focus on the contributions of trajectories that undergo a single act of rescattering. For this orbit subset, one may establish a one-to-one correspondence between the RQSFA and CQSFA contributions for backscattered and forward-scattered trajectory pairs. We assess the influence of the Coulomb potential on the ionization and rescattering times of specific trajectory pairs, kinematic constraints determined by rescattering, and quantum interference between specific pairs of trajectories. We analyze how the Coulomb potential alters their ionization and return times, and their interference in photoelectron momentum distributions. We show that Coulomb effects are not significant for high or medium photoelectron energies and shorter orbits, while, for lower momentum ranges or longer electron excursion times in the continuum, the residual Coulomb potential is more important.
@@ I Introduction
The laser-induced rescattering or recombination of an electron with its parent ion plays a huge role in intense-field laser-matter interaction [1]. This physical picture has been instrumental in explaining a wide range of phenomena, which could not be interpreted by other means, and also in dictating the attosecond s) time scales for which they occurred. These processes may lead to the generation of high-order harmonics [2, 3, 4, 5], or photoelectrons with energies well above the ionization threshold, in what is known as above-threshold ionization (ATI) [6, 7, 8].
@@
The SFA theory was applied to the HATI process in Refs. [34, 26], but the role of the residual Coulomb potential remained an open question. At first, the influence of the Coulomb potential was accounted for through the Born series analyzing the influence of the screening parameter [35], while later approaches incorporate Coulomb distortions in the resulting scattering waves [36], combined with the Born series [37], until eventually the continuum propagation is modified [38, 39, 40, 41], although not all groups of orbits leading to substantial contributions were taken into consideration [38, 39], or approximations have been made on the scattering angle [42, 40, 41]. For a review of the main methods beyond the SFA see the review article [13].
@@
In our earlier publications [50, 53, 54, 58, 46] the comparison between the results obtained using the SFA and the Coulomb quantum-orbit strong-field approximation (CQSFA) was made without considering the structural differences between the theories. For example, in the SFA theory, the intermediate electron momentum changes abruptly at the moment of rescattering which happens when the electron is at the position of the core. In the CQSFA, this is not the case, i.e., the trajectory can go very close to zero, but will never be zero, while the change of the photoelectron momentum is continuous. In the present paper, we present a systematic comparison of the results obtained using our theories. We discuss how the CQSFA trajectories which mimic the behavior of the SFA trajectories can be isolated, and how good is the agreement between the partial contributions to the probability density. Also, we discuss how these differences depend on the photoelectron energy by tracking the phase accumulated due to the Coulomb potential.
The paper is organized as follows. In Section II we present the two versions of the SP method: one based on the SFA and one based on the Coulomb-corrected SFA. In Sec. III we discuss how the trajectories obtained in the SFA theory can be extracted from the CQSFA theory. Moreover, we discuss the behavior of some of the trajectories which do not have an SFA analog. In addition, we briefly present the classification of the quantum orbits for both theories. In Section IV we present our numerical results and discuss the similarities and differences between the two theories. Finally, Section V contains our main conclusions. The atomic units are used unless otherwise stated.
@@ II.2 Coulomb Quantum-Orbit Strong Field Approximation
where and are the integration measures for the path integrals, prime indicates a restriction, and the action reads
@@
A classification of the solutions to the system of equations (11) and (12) was introduced in [50] based upon the initial conditions and the final momentum of the photoelectron trajectory. However, this classification was found to be incomplete [63] due to the fact that more information about the intermediate momentum and position is necessary for a full characterisation of the qualitative nature of the trajectory (see Sec. III). Using the solutions , , and of these Coulomb-corrected SP equations, the probability amplitude can be written as
@@
and is the Maslov phase associated with a solution as calculated by the prescription given in [64] and [65].
@@ III Momentum Mapping and Orbit Correspondence
The backward-scattering solutions are classified using the multiindex . For the return times within one optical cycle, there are infinitely many solutions with different travel times . The index gives the approximate value of the travel time in units of laser-field period , being the angular frequency. For , , there are two pairs of solutions distinguished by the index in such a way that () corresponds to the longer (shorter) travel time. Furthermore, the two solutions of one pair are distinguished by the index . The long (short) solution is denoted with ().
The forward-scattering solutions were classified in [32] using the multiindex in such a way that , measures the travel time, while index is used to distinguish the two solutions of one pair. The positions of the cutoff for these solutions were classically found in [67], while those positions with quantum corrections were given in [68]. Besides the shortest and the second-shortest pair, other forward-scattering solutions contribute only in a low-energy part of the spectra (, being the ponderomotive energy).
For the CQSFA theory, the classification of the solutions was introduced in [50, 54] based upon the initial conditions and the final momentum of the photoelectron trajectory. The solutions with are denoted as ‘orbit 1’ for and ‘orbit 2’ for , while for the solutions are denoted as ‘orbit 3’ for and ‘orbit 4’ for . In those early publications, orbit 1 was identified with a direct SFA orbit in which the electron is freed in the direction of the detector, orbits 2 and 3 behave as field-dressed hyperbolas starting half a cycle earlier or later, with the difference that orbit 2 interacts less with the core and can still be viewed as “direct”, while orbit 3 has a hybrid character. Orbit 4 goes around the core before reaching the detector, and is identified as backscattered. In more recent publications [58, 63], however, we found that these distinctions are blurred and a wide range of behaviors for each of the four types of orbit can be identified by applying spatial filters. If the orbit’s distance of closest approach is larger than the radius determined by the tunnel exit, the CQSFA orbit is considered ‘direct’. If the orbit’s perihelion lies between the tunnel exit and the Bohr radius, that is, , we refer to the orbit as ‘soft scattered’, while if the CQSFA orbit is ‘hard scattered’. A thorough analysis using these filters has been performed in [63].
@@
Using specific subsets of the initial momentum points, rescattered trajectories analogous to those obtained in the SFA can be obtained. In Fig. 1(b) we display these subsets in such a way that the red, blue, orange and pink regions lead to the backward-scattering SFA solutions listed in order of the increasing travel time. For example, the solutions represented by the red points correspond to the backward-scattered SFA solutions with the shortest travel time. Since these solutions can all be classified as orbit 3 or orbit 4, only the solutions classified as such are displayed in Fig.1(b) and Fig.1(c). For the CQSFA we choose the rescattering time to be the time at which the trajectory reaches its perihelion. Physically, this is the most appropriate assumption one can make to compare with the SFA, which is a Born-type expansion for which rescattering times are well defined according to Eq. (6).
The range of these solutions regarding the final momentum is shown in Fig. 1(c), with rescattering ridges marking the maximal classical momenta associated with specific orbits. The rescattering ridges formed by the first- and third-shortest solutions intersect the axis, while this is not the case for the second- and fourth-shortest solutions. The first- and third-shortest solutions with the final momentum (not shown in the figure) correspond to the SFA forward-scattering solutions. Similar solutions do not exist for the second- and fourth-shortest pairs so that the forward-scattering SFA solutions can only be extracted from the orbit 3 CQSFA solutions. The solutions represented by the black dots [see Fig. 1(b)] account for the remaining CQSFA trajectories. These include the forward-scattering solutions and the solutions which include multiple scattering events.
@@
In Fig. 2 we present the values of the initial [panel (a)] and final [panel (b)] photoelectron momentum of the semiclassical orbit 1 trajectories calculated using the CQSFA theory. Among all solutions, we extract those which correspond to the trajectories which make a loop in the clockwise (blue regions) or counterclockwise (red regions) direction around the core. The regions which correspond to the trajectories which do not make a loop around the core are denoted by orange color. The three regions of the final momentum which correspond to these trajectories partiality overlap [see Fig. 2(b)]. However, the ridges of these trajectories can still be easily noticed. The ridge associated to the trajectory which makes a loop in the clockwise direction has the highest energy, while the lowest-energy ridge corresponds to the trajectory which makes a loop in the counterclockwise direction around the core. The examples of these trajectories are shown in Fig. 2(c). The values of the initial-momentum components for the trajectory which makes a loop in the clockwise (counterclockwise) direction are and ( and ), while for the trajectory which does not make a loop around the core, these values are and , in units of . For clarity, although the field dressing means that initially some of the trajectories shown in Fig. 2(c) travel into the 2nd quadrant of the plane, they are indeed orbit 1 trajectories and will eventually move into the 1st quadrant outside of the position ranges shown in the figure. Clearly, these are rescattering trajectories even though they correspond to the orbit 1 solutions. Additionally, the radial vector of the perihelion of the trajectory which loops around the core in a counter-clockwise direction [see the red line in Fig. 2(c)] makes an angle of with the negative axis. This shift of the perihelion is what allows the photoelectron to scatter into the first quadrant and is an effect which relies entirely upon the presence of a long-range Coulomb potential. Similar conclusions are valid for the orbit 2 solutions.
@@ IV.1 Times and orbits
The rescattering events happen around . By comparing the values of the potential during the rescattering event for different values of the photoelectron energy, we see that, for the high and medium values of the energy, the potential rapidly goes to a large negative value (see the blue and orange lines in Fig. 5). On the other hand, for the small value of the photoelectron energy, the behavior of the potential during rescattering is different for the shortest and second-shortest pairs of solutions. In the case of the second-shortest pair, the photoelectron does not closely approach the core and the potential well is approximately as deep as it is at the tunnel exit [see the red lines in Fig. 5(b)]. In addition, the potential exhibits oscillations which continue long after rescattering (see the red lines in Fig. 5). This explains the fact that the deviation of the rescattering time calculated using the CQSFA from the time calculated using the SFA theory is more pronounced in the low-energy region.
Finally, the two branches of the second-shortest (as well as for the fourth-shortest) set of solutions approach each other for low-energy in a similar manner as they do for the high-energy cutoff [see Fig. 4(d)]. This can be related to the results shown in Fig. 1(c). As we already mentioned in Sec. III, the second- and fourth-shortest pairs do not intersect with the axis. Therefore, as the energy of solutions goes to zero in the direction of emission , they approach the caustic and coalesce. This low-energy caustic also passes through the origin in the SFA theory. However, we do not observe the same low-energy coalescence of the rescattering and ionization times. This happens because the presence of a long-range Coulomb potential implies that the photoelectron requires a certain minimum energy to escape this interaction.

### 21 Feb 2024 v2

@@ Abstract
We perform a systematic comparison between photoelectron momentum distributions computed with the rescattered-quantum orbit strong-field approximation (RQSFA) and the Coulomb-quantum orbit strong-field approximation (CQSFA). We exclude direct, hybrid, and multiple scattered CQSFA trajectories, and focus on the contributions of trajectories that undergo a single act of rescattering. For this orbit subset, one may establish a one-to-one correspondence between the RQSFA and CQSFA contributions for backscattered and forward-scattered trajectory pairs. We assess the influence of the Coulomb potential on the ionization and rescattering times of specific trajectory pairs, kinematic constraints determined by rescattering, and quantum interference between specific pairs of trajectories. We analyze how the Coulomb potential alters their ionization and return times, and their interference in photoelectron momentum distributions. We show that Coulomb effects are not significant for high or medium photoelectron energies and shorter orbits, while, for lower momentum ranges or longer electron excursion times in the continuum, the residual Coulomb potential is more important. We also assess the agreement of both theories for different field parameters, and show that it improves with the increase of the wavelength.
@@ I Introduction
The laser-induced rescattering or recombination of an electron with its parent ion plays a huge role in intense-field laser-matter interaction [1, 2]. This physical picture has been instrumental in explaining a wide range of phenomena, which could not be interpreted by other means, and also in dictating the attosecond s) time scales for which they occurred. These processes may lead to the generation of high-order harmonics [3, 4, 5, 6], or photoelectrons with energies well above the ionization threshold, in what is known as above-threshold ionization (ATI) [7, 8, 9].
@@
The SFA theory was applied to the HATI process in Refs. [35, 27], but the role of the residual Coulomb potential remained an open question. At first, the influence of the Coulomb potential was accounted for through the Born series analyzing the influence of the screening parameter [36], while later approaches incorporate Coulomb distortions in the resulting scattering waves [37], combined with the Born series [38]. Eventually, the continuum propagation was modified [39, 40, 41, 42], although not all groups of orbits leading to substantial contributions were taken into consideration [39, 40], or approximations have been made on the scattering angle [43, 41, 42]. For the main methods beyond the SFA see the review article [14].
@@
In our earlier publications [51, 54, 55, 59, 47] the comparison between the results obtained using the SFA and the Coulomb quantum-orbit strong-field approximation (CQSFA) was made without considering the structural differences between the theories. For example, in the SFA theory, the intermediate electron momentum changes abruptly at the moment of rescattering which happens when the electron is at the position of the core. In the CQSFA, this is not the case, i.e., the trajectory can go very close to zero, but will never be zero, while the change of the photoelectron momentum is continuous. Furthermore, we focused on what the SFA leaves out, instead of looking for similarities. The strong-field approximation was presented as a limit for the CQSFA orbits in specific parameter ranges [59], or comparisons with direct SFA electrons were performed to emphasize the deflection caused by the residual potential [55, 47]. Up to recently, the CQSFA was solved as a purely boundary problem, which required some pre-knowledge of the orbits’ dynamics. This excluded entire sets of orbits, which were revealed in a recent publication in which hybrid and forward implementations of the CQSFA were made [60]. These orbits led to annular intra-cycle fringes that resembled those found for rescattered SFA orbits [61].
In the present paper, we systematically compare the results obtained using our theories. We discuss how the CQSFA trajectories which mimic the behavior of the SFA trajectories can be isolated, and how good is the agreement between the partial contributions to the probability density. Also, we discuss how these differences depend on the photoelectron energy by tracking the phase accumulated due to the Coulomb potential. Our focus will be on the rescattered regime, as a comparison with direct SFA electrons has been performed elsewhere [54, 55].
To keep the focus on how far a Born-type approach goes in reproducing the full continuum dynamics, we have opted for the simplest possible system: hydrogen in a linearly polarized monochromatic field. Simulating a real experiment may require focal averaging, depletion, bound-state dynamics, and a realistic pulse. These effects are potential sources of incoherence [62], and they may skew the interference patterns one is trying to assess (see [63, 64] for examples in a two-electron scenario), and modify the dynamical constraints associated with rescattering. Assessing all these additional effects would detract from the main purpose of the present work. However, it is known that for pulses longer than 10 optical cycles a flat envelope is a good approximation [65]. Furthermore, the CQSFA has been successfully employed to explain experiments in rare gases such as xenon [46] and argon [66, 67], as well as to detect parity in diatomic molecules [68]. In the modeling of these experiments, a monochromatic wave was assumed, and, still, the theory reproduced subtle effects such as multi-path holographic interference [66, 67].
The paper is organized as follows. In Section II we present the two versions of the saddle-point method: one based on the SFA and one based on the Coulomb-corrected SFA. In Sec. III we discuss how the trajectories obtained in the SFA theory can be extracted from the CQSFA theory. Moreover, we discuss the behavior of some of the trajectories which do not have an SFA analog. In addition, we briefly present the classification of the quantum orbits for both theories. In Section IV we present our numerical results and discuss the similarities and differences between the two theories. Finally, Section V contains our main conclusions. The atomic units are used unless otherwise stated.
@@ II.2 Coulomb Quantum-Orbit Strong Field Approximation
where and are the integration measures for the path integrals, the prime indicates a restriction, and the action reads
@@
The integral and the saddle-point equations are solved using a two-pronged contour. The first part of the contour is chosen along the imaginary-time axis, from to , and the second part of the contour is taken along the real-time axis from to a final time . In the first part of the contour, the momentum is assumed to be constant and the binding potential in Eq. (11) is neglected. This makes the sub-barrier contribution to the action dependent on the tunnel trajectory.
which can be used to define the tunnel exit
where is the component of the tunnel trajectory [Eq. (13) for ] parallel to the driving-field polarization. One should note that setting the tunnel exit to be real is an approximation, which renders the continuum propagation equations real. This approximation leads to practically no changes in the interference patterns and has been used in almost all our publications. Solving the full complex problem will involve branch cuts and has been discussed elsewhere [73].
A classification of the solutions to the system of equations (11) and (12) was introduced in [51] based upon the initial conditions and the final momentum of the photoelectron trajectory. However, this classification was found to be incomplete [60] because more information about the intermediate momentum and position is necessary for a full characterization of the qualitative nature of the trajectory (see Sec. III). Using the solutions , , and of these Coulomb-corrected SP equations, the probability amplitude can be written as
@@
, and is the Maslov phase associated with a solution as calculated by the prescription given in [74] and [75]. Here, we use the hybrid CQSFA implementation discussed in [60].
@@ III.1 Orbit Classification
The backward-scattering solutions are classified using the multiindex . For the return times within one optical cycle, there are infinitely many solutions with different travel times . The index gives the approximate value of the travel time in units of laser-field period , being the angular frequency. For , , there are two pairs of solutions distinguished by the index in such a way that () corresponds to the longer (shorter) travel time. Furthermore, the two solutions of one pair are distinguished by the index . The long (short) solution is denoted by ().
The forward-scattering solutions were classified in [33] using the multiindex in such a way that , measures the travel time, while index is used to distinguish the two solutions of one pair. The positions of the cutoff for these solutions were classically found in [77], while those positions with quantum corrections were given in [78]. Besides the shortest and the second-shortest pair, other forward-scattering solutions contribute only in a low-energy part of the spectra [, being the ponderomotive energy].
For the CQSFA theory, the classification of the solutions was introduced in [51, 55] based upon the initial conditions and the final momentum of the photoelectron trajectory. The solutions with are denoted as ‘orbit 1’ for and ‘orbit 2’ for , while for the solutions are denoted as ‘orbit 3’ for and ‘orbit 4’ for , where is the tunnel exit. In those early publications, orbit 1 was identified with a direct SFA orbit in which the electron is freed in the direction of the detector, orbits 2 and 3 behave as field-dressed hyperbolas starting half a cycle earlier or later, with the difference that orbit 2 interacts less with the core and can still be viewed as “direct”, while orbit 3 has a hybrid character. Orbit 4 goes around the core before reaching the detector, and is identified as backscattered. In more recent publications [59, 60], however, we found that these distinctions are blurred and a wide range of behaviors for each of the four types of orbit can be identified by applying spatial filters. If the orbit’s distance of closest approach is larger than the radius determined by the tunnel exit, the CQSFA orbit is considered ‘direct’. If the orbit’s perihelion lies between the tunnel exit and the Bohr radius, that is, , we refer to the orbit as ‘soft scattered’, while if the CQSFA orbit is ‘hard scattered’. A thorough analysis using these filters has been performed in [60].
@@ III.2 Extracting SFA solutions
Using specific subsets of the initial momentum points, rescattered trajectories analogous to those obtained in the SFA can be obtained. In Fig. 1(b) we display these subsets in such a way that the red (gray), blue (dark gray), orange (light gray) and pink (bright gray) regions lead to the backward-scattering SFA solutions listed in order of the increasing travel time. For example, the solutions represented by the red (gray) points correspond to the backward-scattered SFA solutions with the shortest travel time. Since these solutions can all be classified as orbit 3 or orbit 4, only the solutions classified as such are displayed in Fig.1(b) and Fig.1(c). For the CQSFA we choose the rescattering time to be the time at which the trajectory reaches its perihelion. Physically, this is the most appropriate assumption one can make to compare with the SFA, which is a Born-type expansion for which rescattering times are well defined according to Eq. (6).
The range of these solutions regarding the final momentum is shown in Fig. 1(c), with rescattering ridges marking the maximal classical momenta associated with specific orbits. The rescattering ridges formed by the first- and third-shortest solutions intersect the axis, while this is not the case for the second- and fourth-shortest solutions. The first- and third-shortest solutions with the final momentum (not shown in the figure) correspond to the SFA forward-scattering solutions. Similar solutions do not exist for the second- and fourth-shortest pairs so that the forward-scattering SFA solutions can only be extracted from the orbit 3 CQSFA solutions. The solutions represented by the black dots [see Fig. 1(b)] account for the remaining CQSFA trajectories. These include the forward-scattering solutions and the solutions which include multiple scattering events. A detailed study of the initial-to-final momentum mapping of the CQSFA orbits, with and without SFA counterparts, has been performed in [60], using different versions of the CQSFA and the spatial filtering mentioned in Sec. III.1.
III.3 Counterexamples: orbits with no SFA analogy
@@
In Fig. 2 we present the values of the initial [panel (a)] and final [panel (b)] photoelectron momentum of the semiclassical orbit 1 trajectories calculated using the CQSFA theory. Among all solutions, we extract those which correspond to the trajectories which make a loop in the clockwise [blue (dark gray) regions] or counterclockwise [red (gray) regions] direction around the core. The regions which correspond to the trajectories which do not make a loop around the core are denoted by orange (light gray) color. The three regions of the final momentum which correspond to these trajectories partiality overlap [see Fig. 2(b)]. However, the ridges of these trajectories can still be easily noticed. The ridge associated to the trajectory which makes a loop in the clockwise direction has the highest energy, while the lowest-energy ridge corresponds to the trajectory which makes a loop in the counterclockwise direction around the core. The examples of these trajectories are shown in Fig. 2(c). The values of the initial-momentum components for the trajectory which makes a loop in the clockwise (counterclockwise) direction are and ( and ), while for the trajectory which does not make a loop around the core, these values are and , in units of . For clarity, although the field dressing means that initially some of the trajectories shown in Fig. 2(c) travel into the 2nd quadrant of the plane, they are indeed orbit 1 trajectories and will eventually move into the 1st quadrant outside of the position ranges shown in the figure. Clearly, these are rescattering trajectories even though they correspond to the orbit 1 solutions. Additionally, the radial vector of the perihelion of the trajectory which loops around the core in a counter-clockwise direction [see the red (gray) line in Fig. 2(c)] makes an angle of with the negative axis. This shift of the perihelion is what allows the photoelectron to scatter into the first quadrant and is an effect which relies entirely upon the presence of a long-range Coulomb potential. Similar conclusions are valid for the orbit 2 solutions.
The examples discussed above focus on the specific scenarios for which there is hard scattering, in the sense that the orbit’s distance of closest approach is smaller than that defined by the tunnel exit. However, in previous publications we have provided examples of many orbits which do not have SFA counterparts. For instance, in [54] we have shown that the near-threshold fan-shaped structure encountered in experiments occurs due to a Coulomb distortion of orbits 2, which is angle-dependent and resembles a laser-dressed Kepler hyperbola. Upon interference with orbits 1, it gives rise to a radial set of fringes centered at . A direct orbit 2 in the SFA sense, instead of the CQSFA orbit, leads to nearly vertical structures, and fails to reproduce the fan [55]. Furthermore, orbit 3 cannot always be understood as “forward scattered”, and under some circumstances exhibits a hybrid character. This is important for an accurate description of the spider-like fringes [55, 59]. In [59], we reported several overlooked holographic structures, whose existence is determined by the existence of the residual potential in the continuum. One of these structures, the spiral, has been identified in experiments [46], and has been shown not to be reproduced by either direct or rescattered ATI using the standard SFA.
@@ IV.1 Times and orbits
The rescattering events happen around . By comparing the values of the potential during the rescattering event for different values of the photoelectron energy, we see that, for the high and medium values of the energy, the potential rapidly goes to a large negative value (see the blue (dark gray) and orange (light gray) lines in Fig. 5). On the other hand, for the small value of the photoelectron energy, the behavior of the potential during rescattering is different for the shortest and second-shortest pairs of solutions. In the case of the second-shortest pair, the photoelectron does not closely approach the core and the potential well is approximately as deep as it is at the tunnel exit [see the red (gray) lines in Fig. 5(b)]. In addition, the potential exhibits oscillations which continue long after rescattering (see the red (gray) lines in Fig. 5). This explains the fact that the deviation of the rescattering time calculated using the CQSFA from the time calculated using the SFA theory is more pronounced in the low-energy region.
Finally, the two branches of the second-shortest (as well as for the fourth-shortest) set of solutions approach each other for low-energy in a similar manner as they do for the high-energy cutoff [see Fig. 4(d)]. This can be related to the results shown in Fig. 1(c). As we already mentioned in Sec. III, the sets of saddle point solutions corresponding to the second- and fourth-shortest pairs do not intersect with the axis. Therefore, as the energy of solutions goes to zero in the direction of emission , they approach the caustic and coalesce. This low-energy caustic also passes through the origin in the SFA theory and, equivalently, a coalescence of saddles occurs. However, we do not observe the same low-energy coalescence of the rescattering and ionization times since the coalescence is only visible on extremely low energy scales. This happens because the presence of a long-range Coulomb potential implies that the photoelectron requires a certain minimum energy to escape this interaction.
It is also worth noting that, although the rescattering and ionization times of the third-shortest pair of solutions begin to approach for low energies, they do not actually coalesce. The coalescence occurs, as is the case for the shortest pair, in the half plane with opposite sign of . This is the momentum region occupied by the solutions analogous to the SFA forward scattered solutions. It has been excluded by the restrictions imposed in the CQSFA to match its orbits with the backscattered SFA solutions. For the shortest pair, this region will give rise to the solution, and for the third-shortest pair it corresponds to the forward-scattered solutions. There are no forward-scattering orbits corresponding to the remaining pairs. The beginning of this coalescence is visible for the third-shortest pair of solutions because the intersection of the set of saddle point solutions for the third-shortest pair intersects with the axis at a much smaller perpendicular momentum than the set of saddle-point solutions for the shortest pair.